An express train makes a run of 240 km at a certain speed.Another train whose speed is 12 km/hr less takes an hour longer to cuver the same distance. Find the speed of the express train in km/hr.
Answer by lwsshak3(11628)

An express train makes a run of 240 km at a certain speed.Another train whose speed is 12 km/hr less takes an hour longer to cuver the same distance. Find the speed of the express train in km/hr.
**
let x=speed of express train
x-12=speed of slower train
travel time=distance/speed
..
240/(x-12)-240/x=1
LCD: x(x-12)
240x-240x+2880=x^2-12x
x^2-12x-2880
(x+48)(x-60)=0
x=-48(reject x>0)
or
x=60
ans:
Speed of the express train=60 km/hr